OpenStudy (anonymous):

Harry had $32. He spent all the money on buying 3 notebooks for $x each and 4 packs of index cards for $y each. If Harry had bought 5 notebooks and 5 packs of index cards, he would have run short of $18. A student concluded that the price of each notebook is $5 and the price of each pack of index cards is $1. Which statement best justifies whether the student's conclusion is correct or incorrect?

OpenStudy (anonymous):

The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 50 is (8, 2). The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (8, 2). The student's conclusion is correct because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (5, 1). The student's conclusion is correct because the solution to the system of equations 3x – 4y = 32 and 5x – 5y = 50 is (5, 1).

OpenStudy (anonymous):

Let x be cost notebook and y be cost of index card According to first condition 3x+4y=32 According to second comdition 5x+5y=32+15 Solve the two equations

OpenStudy (anonymous):

x=4 and y=5?

OpenStudy (anonymous):

@tanmaykhandait

OpenStudy (anonymous):

No...my error... In the second equation.....it is 5x+5y=32+18

OpenStudy (anonymous):

Really sorry

OpenStudy (anonymous):

Y=2 and x=8

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

@tanmaykhandait so A,right?

OpenStudy (anonymous):

Yup